OpenStudy (anonymous):

Can Someone Help me with this? Find the equation of the line that passes through the points (-2, 1) and (1, -2).

OpenStudy (anonymous):

This can be found by using two point formula: \[\large \frac{y - y_1}{y_2 - y_1} = \frac{x-x_1}{x_2 - x_1}\]

OpenStudy (anonymous):

\[y-1\div1-(-2)=x-(-2)\div1-(-2)\]

OpenStudy (anonymous):

@waterineyes am Iin right way?

OpenStudy (anonymous):

denominator is wrong.. It should be -2 -1 .. Check it according to the formula I have written above..

OpenStudy (anonymous):

sorry I always mixed up my eyes! @waterineyes

OpenStudy (anonymous):

For left hand side.. Right hand side it is right..

OpenStudy (anonymous):

\[\frac{y-1}{-2 - 1} = \frac{x- (-2)}{1 - (-2)}\]

OpenStudy (anonymous):

\[\frac{y-1}{-3} = \frac{x +2}{3} \implies y - 1 = -x + 2\] \[\large \color{green}{x + y = 3}\]

OpenStudy (anonymous):

y-1/-3=x+2/3

OpenStudy (anonymous):

Yes carry on..

OpenStudy (anonymous):

Cross multiply now..

OpenStudy (anonymous):

do I need to do cross multiplication after that?

OpenStudy (anonymous):

3y-3=-3x+6?

OpenStudy (anonymous):

or you can multiply -3 both the sides.. Do this one this will be easy for you..

OpenStudy (anonymous):

y+3=-3?

OpenStudy (anonymous):

y+x=-3 I mean @waterineyes

OpenStudy (anonymous):

Again solve.. try it once more.. are you applying here cross multiplication??

OpenStudy (anonymous):

y+x=3 double check it

OpenStudy (anonymous):

Yes it is right.. Well Done..

OpenStudy (anonymous):

Thanks @waterineyes

OpenStudy (anonymous):

Welcome dear..